OpenStudy (anonymous):

The scientific notation 4.87 × 10-3 has what value?

OpenStudy (mertsj):

.00487

OpenStudy (lollylau):

10^-3 = 0.001. 4.87 * 10^-3 =4.87 * 0.001 = 0.00487

OpenStudy (amistre64):

a neg. exponent is small, so we have to make the number part smaller